OpenStudy (anonymous):

Simplify the rational expression: 3x-15/3x+24

Directrix (directrix):

(3x-15)/(3x+24) = 3*(x -5) / 3*(x + 8) = ? after the threes are divided out?

Directrix (directrix):

@severalgoats